OpenStudy (anonymous):

I need some help setting this problem up to find all the solutions to this equation. sec(θ/2) = cos (θ/2)

OpenStudy (anonymous):

Do I use the half angle formula? And if so, how do I handle the sec(θ/2)? I know sec is 1/cos but I'm not sure how to factor that into the formula...

jimthompson5910 (jim_thompson5910):

since sec(theta) = 1/cos(theta) we go from this sec(θ/2) = cos (θ/2) to this 1/cos(θ/2) = cos (θ/2)

jimthompson5910 (jim_thompson5910):

then you multiply both sides by cos(θ/2) 1/cos(θ/2) = cos (θ/2) 1 = cos (θ/2)*cos(θ/2) 1 = cos^2 (θ/2) cos^2 (θ/2) = 1 do you see what to do next?

OpenStudy (anonymous):

When I multiply both sides by cos(θ/2) I get θ/2 = cos^2(θ/2) ?

jimthompson5910 (jim_thompson5910):

think of cos(θ/2) as x

jimthompson5910 (jim_thompson5910):

so 1/cos(θ/2) = cos (θ/2) is really the same as 1/x = x

OpenStudy (anonymous):

Ok, let me write this out again on my paper...

OpenStudy (jdoe0001):

\(\bf sec\left(\frac{\theta}{2}\right)=cos\left(\frac{\theta}{2}\right)\implies \cfrac{1}{cos\left(\frac{\theta}{2}\right)}=cos\left(\frac{\theta}{2}\right) \\ \quad \\ 1=cos^2\left(\frac{\theta}{2}\right)\implies \sqrt{1}=\sqrt{cos^2\left(\frac{\theta}{2}\right)}\implies 1=cos\left(\frac{\theta}{2}\right) \\ \quad \\ 1=\pm\sqrt{\cfrac{1+cos(\theta)}{2}}\implies 1^2=\pm\left[\sqrt{\cfrac{1+cos(\theta)}{2}}\right]^2\implies 1=\pm\cfrac{1+cos(\theta)}{2}\)

OpenStudy (jdoe0001):

as pointed out by jim_thompson5910 , so you'd end up using the half-angle identity

OpenStudy (anonymous):

The way you wrote out 1/cos(θ/2) was helpful. I know it's the same thing, but seeing it written that way makes it make sense.

OpenStudy (anonymous):

I know I'm slow, but I think I figured it out. cos(θ) must then equal 1, which is 2π. So my answers would then be 2π + 2πk.

jimthompson5910 (jim_thompson5910):

and you can simplify that to say x = 2πk where k is any integer

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome